AQ2 Data Units [x] Cost [C] Rev [R] Profit [P] 100 260000 47500 -212500 250 275000...

AQ2
Data Units [x] Cost [C] Rev [R] Profit [P]
100 260000 47500 -212500
250 275000 109375 -165625
500 300000 187500 -112500
800 330000 240000 -90000
1000 350000 250000 -100000
1200 370000 240000 -130000
1300 380000 227500 -152500

As your Division’s Chief economist, you perform a periodic review of the current total Cost "C", Revenue "R" and Profit "P" models associated with one of the Division's newest, but possibly underperforming, product lines. Among other important questions, you would like answers to the ones provided below. [the dependent variables are “C”, “R” or “P”; the independentvariable is “x”, the number of units (produced and) sold]. The "best fit" linear or quadraticC”, “R” or “P” models, derived from the data referenced above, follow:

    C = 100(x)1 + 250,000 with [(R^2)] = 1.0.

    R = –0.25(x)2 + 500(x)1 with [(R^2)] = 1.0.

    P = –0.25(x)2 + 400(x)1 – 250,000 with [(R^2)] = 1.0.  

1. Calculate how many product units “x” should be sold per sales period to maximizethis product’s total Profit “P”…then…calculate “Pmax” at this “x” value. Assume market constraints suggest that 1,400 is the maximum number of product units that actually can be sold per sales period (i.e. 0 < x < 1,400 units).  

2. If internal production or external market constraints are not an issue, determine, if possible, the minimum number of product units “x” that must be produced and sold in order to “break even” (i.e. generate a Profit of $0.00)?
Answer #1

1. The profit P = -0.25x^2 + 400x - 250000 would be maximum where

\frac{\mathrm{d} P}{\mathrm{d} x} = 0

or \frac{\mathrm{d} }{\mathrm{d} x}(-0.25x^2 + 400x - 250000) = 0

or -0.5x + 400 = 0

or 0.5x = 400

or x = 800 units.

Hence, profit would be maximum when production level is 800 units, which satisfies the constraint 0<x<1400.

The maximum profit would be as P_{max} = -0.25*800^2 + 400*800 - 250000 or P_{max} = -90000 (dollars). The negative profit indicates that the loss would be minimum at x=800, and the loss would be 90000.

2. The profit when maximum is negative, which implies that the profit never reaches $0. The maximum profit is -$90000, and since it is non-positive, the profit is never zero. Also, we have the the average cost as AC = \frac{C}{x} = 100 + \frac{250000}{x} , which is asymptotic to the AC=100 horizontal line. Hence, theoretically, AC would be minimum where x = \infty , and minimum AC would be 100 in that case.

As the profit never reaches zero, and since AC is minimum for a very large unattainable amount of x, it is not possible to determine the breakeven amount of x.